# Adversarially Perturbed Precision Matrix Estimation

### v1

@@ 3.1.1 Properties
Corollary 3.2 (Sparsity Recovery).
Let $\widetilde{C}$ be the minimizer of problem (9). Suppose that $\delta=\eta/\sqrt{n}$ with $\eta>0$. Then, as $n\to\infty$, the asymptotic distribution of $\widetilde{C}]_{ij}$ has a positive probability mass at 0 whenever $\Sigma^{-1}]_{ij}=0$.
Remark 3.3.
Corollary 3.2 follows immediately from Theorem 5.2 in Section 5 and Proposition 3.8 in Xie and Huo (2025).
Corollary 3.2 implies that for zero entries in the true precision matrix $\Sigma^{-1}$, the asymptotic distribution of the corresponding estimates from (9) has a positive probability at zero. We call this property sparsity recovery, which is theoretically guaranteed if the adversarial perturbation order is $1/\sqrt{n}$. The sparsity is induced by the first penalty term in (9). Notably, while similar to the shrinkage achieved in $\ell_{1}$-regularized precision matrix estimation (Yuan and Lin, 2007; Friedman et al., 2008), this term incorporates the adaptive adjustment for adversarial robustness discussed earlier
@@ 4 Regularization Effect
We characterize this regularization effect mathematically in the following proposition.
@@
Suppose the random variable $\bm{x}$ follows a distribution $\bm{P}$ with $\mathbb{E}_{\bm{P}}[\|\bm{x}\|_{p}]<\infty$ and $\bm{P}([C\bm{x}]_{i}=0)=0$, $C$ is a positive definite matrix with $\|C\|_{p\to q}<\infty$, and $p\in(1,\infty]$ with $1/p+1/q=1$. Then, we have that
@@
$\bm{v}_{C\bm{x}}=\frac{\operatorname{sign}(C\bm{x})\odot|C\bm{x}|^{q-1}}{\|C\bm{x}\|_{q}^{q-1}}.$ (10)
@@
Since the term $\log\det C$ in the proposed framework (2) is deterministic, Proposition 4.1 focuses on the inner maximization problem in (2) to analyze the induced regularization under $\ell_{p}$-perturbation. Instead of only analyzing on the empirical distribution, we consider a more general distribution $\bm{P}$ satisfying certain regularity conditions. The stated regularity conditions in Proposition 4.1 ensure well-posedness and integrability.
@@
$=$ $\mathbb{E}_{\bm{P}}[\bm{x}^{\top}C\bm{x}]+2\delta\mathbb{E}_{\bm{P}}\left[\|C\bm{x}\|_{1}\right]+\delta^{2}\mathbb{E}_{\bm{P}}\left[\operatorname{sign}(C\bm{x})^{\top}C\operatorname{sign}(C\bm{x})\right]+o(\delta^{2}),$
@@
The first-order regularization term arises from Hölder’s inequality and penalizes the $\ell_{q}$ norm of the gradient direction, where $1/p+1/q=1$. The vector $\bm{v}_{C\bm{x}}$ in the second-order term is the $\ell_{p}$-unit vector that achieves the associated Hölder dual pairing, as defined in (10). The second-order regularization term reflects the local curvature of the worst-case loss. Specifically, this term corresponds to the mean of the Rayleigh quotient of $C$ evaluated with respect to the vector $C\bm{x}$ (when $p=2$) or $\operatorname{sign}(C\bm{x})$ (when $p=\infty$). If $C$ has the eigen-decomposition $C=\sum_{j}\lambda_{j}\bm{u}_{j}\bm{u}_{j}^{\top}$, the associated Rayleigh quotient represents a weighted average of the eigenvalues. For $p=2$, it takes the form:
@@
$\operatorname{sign}(C\bm{x})^{\top}C\operatorname{sign}(C\bm{x})=\sum_{j}\lambda_{j}(\bm{u}_{j}^{\top}\operatorname{sign}(C\bm{x}))^{2}.$
@@ 5 Asymptotic Distribution
$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}([\bm{x}]_{i}[\bm{x}]_{j},[\bm{x}]_{i^{\prime}}[\bm{x}]_{j^{\prime}}).$
@@
Firstly, the proposed perturbed estimation $\widehat{C}$ is a consistent estimator of the precision matrix $\Sigma^{-1}$, i.e., $\widehat{C}\to_{p}\Sigma^{-1}$, when the perturbation magnitude shrinks as the sample size increases. Then, the proposed perturbed estimation has a convergence order of $\sqrt{n}$ when $\gamma\geq 1/2$ and a slower convergence order of $n^{\gamma}$ when $\gamma\in(0,1/2)$.
@@
$\widetilde{C}\in\arg\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\omega_{k}|[C]_{kj}|+\delta^{2}\|C\|_{1,1}\right\}.$ (13)
@@
    $+\sqrt{\frac{8}{\pi}}\eta\sum_{j=1}^{d}\sum_{k=1}^{d}[\Sigma]_{kk}\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)\right$.
    $\left.+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
@@
    $+\sqrt{\frac{8}{\pi}}\eta\sum_{j=1}^{d}\sum_{k=1}^{d}[\Sigma]_{kk}\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)\right$.
    $\left.+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
@@
$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}([\bm{x}]_{i}[\bm{x}]_{j},[\bm{x}]_{i^{\prime}}[\bm{x}]_{j^{\prime}}).$
@@
Theorem 5.2 provides the theoretical basis for establishing the sparsity recovery property of the estimation $\widetilde{C}$, as formalized in Corollary 3.2
Theorem 5.2 shows that the estimation $\widetilde{C}$ in problem (13) shares the same asymptotic normal distribution of the estimation $\widehat{C}$ in problem (12). However, for $\gamma=1/2$, the asymptotic distribution of $\widetilde{C}$ becomes non-normal. Similar to the standard $\ell_{1}$-regularized precision matrix estimation (Yuan and Lin, 2007), this non-normality helps encourage sparsity in some sense. Note that the asymptotic term $\Sigma]_{kk}$ presented in Theorem 5.2 represents the impact of adversarial perturbation in our surrogate framework (13).
@@
$\in(1,\infty] p=p$$\infty$, surrogate Rate
@@ 6 Numerical Experiments
In this section, we will deliver numerical experiments for the proposed adversarially perturbed precision matrix estimation scheme and compare it with the $\ell_{1}$-regularized precision matrix estimation (Friedman et al., 2008). The experiments are implemented by R.
6.1 Computation
We will focus on the surrogate $\ell_{\infty}$-perturbed matrix estimation framework (9) discussed in Section 3.
Due to the symmetry of the matrix $C$, we can reformulate problem (9) equivalently as follows,
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathrm{tr}(\bar{A}C)+\|C\|_{1,\Lambda}\right\}$ (14)
where
$\bar{A}=\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}],\quad\|C\|_{1,\Lambda}=\sum_{k=1}^{d}\sum_{j=1}^{d}[\lambda]_{kj}\left|[C]_{kj}\right|$
$\lambda]_{kj}=\delta\mathbb{E}_{\mathbb{P}_{n}}\left[\left|[\bm{x}]_{k}\right|\right]+\delta\mathbb{E}_{\mathbb{P}_{n}}\left[\left|[\bm{x}]_{j}\right|\right]+\delta^{2}.$ (15)
The notation $\|C\|_{1,\Lambda}$ denotes the weighted element-wise $\ell_{1}$-norm of the matrix $C$ with the symmetric weight matrix $\Lambda$. The components of the weight matrix $\Lambda$ are dependent on the moment of the observations and the perturbation magnitude as shown in (15). In this way, problem (14) can be considered as the generalized sparse inverse covariance estimation problem discussed in Hsieh et al. (2014). Thus, problem (14) can be solved by the QUIC algorithm proposed by Hsieh et al. (2014). The QUIC algorithm is a second-order algorithm and has been proven to be superlinearly convergent (Hsieh et al., 2014).
In addition, immediately from Theorem 9 in Hsieh et al. (2014), we can conclude that the precision matrix estimation obtained from the $\ell_{\infty}$-perturbed precision matrix estimation proposed in Section 3 is unique.
Corollary 6.1.
There exists a unique minimizer $\widetilde{C}$ for the optimization problem (14) when $\delta>0$.
As the $\ell_{\infty}$-perturbed precision matrix estimation has the weight matrix dependent on the observed samples, the $\ell_{1}$-regularized precision matrix estimation (Friedman et al., 2008) has a constant weight, as shown in the following:
@@ 6.1 Computation
We also apply the QUIC algorithm (Hsieh et al., 2014) to solve the optimization problem above.
@@
6.2 Experiments with Synthetic Data
In this subsection, we will carry out the numerical experiments on the following synthetic models, as suggested in Yuan and Lin (2007).
Model 1. An $AR(1)$ model with $c_{i,i}=1$ and $c_{i,i-1}=c_{i-1,i}=0.5$, for i=1,,di=1,...,d.
Model 2. An $AR(2)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.5$ and $c_{i,i-2}=c_{i-2,i}=0.25$, for i=1,,di=1,...,d.
Model 3. An $AR(3)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.4$ and $c_{i,i-2}=c_{i-2,i}=c_{i,i-3}=c_{i-3,i}=0.2,$ for i=1,,di=1,...,d.
Model 4. An $AR(4)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.4$, $c_{i,i-2}=c_{i-2,i}=c_{i,i-3}=c_{i-3,i}=0.2$ and $c_{i,i-4}=c_{i-4,i}=0.1,$ for i=1,,di=1,...,d.
Model 5. Star model with every node connected to the first node, with $c_{i,i}=1$, $c_{1,i}=c_{i,1}=0.2$ and $c_{i,j}=0$ otherwise, for i,j=1,,di,j=1,...,d.
Model 6. Circle model with $c_{ii}=1$, $c_{i,i-1}=c_{i-1,i}=0.5$ and $c_{1,d}=c_{d,1}=0.4$, for i=1,,di=1,...,d.
We may use the following metrics to evaluate the numerical performance in the rest of this section: true negative rate (TNR), true positive rate (TPR), accuracy (ACC), and Matthews Correlation Coefficient (MCC) (Guilford, 1954). The metrics are defined as follows:
@@ 6.2 Experiments with Synthetic Data
$\mathrm{ACC}=\frac{\mathrm{TN}+\mathrm{TP}}{\mathrm{TN}+\mathrm{FP}+\mathrm{TP}+\mathrm{FN}}$

### v2

@@ 3.1 Adversarial Robustness
3.2 High-dimensional Consistency
This subsection will investigate the high-dimensional convergence rate and model selection consistency for the estimation obtained from (9). The analysis is based on the primal-dual witness technique innovated by wainwright2009sharp. Our analysis extends the primal-dual witness technique to scenarios involving data-dependent and non-uniform penalty parameters.
We first reformulate problem (9) equivalently as follows
$\widetilde{C}\in\arg\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}^{\top}C\bm{x}]+\sum_{k\neq j}\hat{\lambda}_{kj}\left|C_{kj}\right|\right\},$ (10)
$\hat{\lambda}_{kj}=\delta\mathbb{E}_{\mathbb{P}_{n}}\left[\left|\bm{x}_{k}\right|\right]+\delta\mathbb{E}_{\mathbb{P}_{n}}\left[\left|\bm{x}_{j}\right|\right]+\delta^{2},$
where we do not penalize the diagonals to simplify the analysis.
We then present several notations and necessary assumptions for our subsequent analysis.
Let the off-diagonal support index set of the ground-truth precision matrix $\Sigma^{-1}$ be $E:=\{(i,j):i\not=j,[\Sigma^{-1}]_{ij}\neq 0,i,j=1,\dots,d\},$ and the augmented support index set with diagonals included is denoted by $S:=E\cup\{(i,i):i=1,\dots,d\}.$ Define the maximum number of nonzeros per row in $\Sigma^{-1}$ with diagonals counted as $s:=\max_{1\leq i\leq d}\#\{j:\ [\Sigma^{-1}]_{ij}\neq 0,j=1,\dots,d\}.$ Let $\omega_{i}^{\ast}:=\mathbb{E}\left[\left|\bm{x}_{i}\right|\right]=\sqrt{2{(\pi)^{-1}}\Sigma_{ii}}$ and define $\lambda^{\ast}_{ij}:=\delta(\omega_{i}^{\ast}+\omega_{j}^{\ast})+\delta^{2},i\neq j,\lambda^{\ast}_{ii}:=0,i,j=1,\dots,d$. Let $\Lambda^{\ast}$ be the symmetric matrix with entries $\lambda^{\ast}_{ij}$. Let $\sigma_{\max}:=\max_{1\leq i\leq d}\sqrt{\Sigma_{ii}}$, $\sigma_{\min}:=\min_{1\leq i\leq d}\sqrt{\Sigma_{ii}}$. Define $\omega_{\max}^{\ast}:=\max_{1\leq i\leq d}\omega_{i}^{\ast}=\sqrt{2{(\pi)^{-1}}}\sigma_{\max},$ $\omega_{\min}^{\ast}:=\min_{1\leq i\leq d}\omega_{i}^{\ast}=\sqrt{2{(\pi)^{-1}}}\sigma_{\min}.$ We also let Hessian take the form $\Gamma=\Sigma\otimes\Sigma$.
Assumption 3.2 (Covariance Control).
$\kappa_{\Gamma}:=\|(\Gamma_{SS})^{-1}\|_{\infty}<\infty,\qquad\kappa_{\Sigma}:=\|\Sigma\|_{\infty}<\infty.$
Assumption 3.3 (Scale-adaptive Incoherence Condition).
Define $A:=\Gamma_{S^{c}S}(\Gamma_{SS})^{-1},$ and
$\kappa_{A}:=\|A\|_{\infty}<\infty,\quad\psi^{\ast}:=\max_{e\in S^{c}}\sum_{s\in S}|A_{es}|\frac{\lambda^{\ast}_{s}}{\lambda^{\ast}_{e}}\leq 1-\alpha,\quad\alpha\in(0,1).$ (11)
Theorem 3.4 (Convergence Rate and Model Selection Consistency).
Suppose Assumptions 3.23.3 hold. Fix any $\tau>2$ and set
$\delta=c_{\delta}\sqrt{\frac{\log d}{n}},\qquad c_{\delta}=\frac{\sqrt{\tau}}{\alpha}\max\left\{16C_{2}\sigma_{\max},\ \ \frac{2\sqrt{2\pi}(1+\kappa_{A})C_{1}\sigma_{\max}^{2}}{\sigma_{\min}}\right\},$
where $C_{1},C_{2}$ are some universal constants. Define the constant $B:=C_{1}\sigma_{\max}^{2}\sqrt{\tau}+3\sqrt{2{(\pi)^{-1}}}\sigma_{\max}c_{\delta}+c_{\delta}^{2}.$ Assume the sample size satisfies
$n\geq\max\left\{2\pi C_{2}^{2}\left(\frac{\sigma_{\max}}{\sigma_{\min}}\right)^{2}\tau,36\kappa_{\Gamma}^{4}\kappa_{\Sigma}^{6}B^{2}s^{2},\left(\frac{12\sqrt{2\pi}(1+\kappa_{A})\kappa_{\Sigma}^{3}\kappa_{\Gamma}^{2}B^{2}}{\alpha\sigma_{\min}c_{\delta}}\right)^{2}s^{2}\right\}\log d.$
Then, for the solution $\widetilde{C}$ to problem (10), with probability at least $1-4d^{2-\tau}$, the following holds:
$\|\widetilde{C}-\Sigma^{-1}\|_{\max}\ \leq\ 2\kappa_{\Gamma}B\sqrt{\frac{\log d}{n}},\quad\widetilde{C}_{S^{c}}=0.$
If the inequality
$\min_{(i,j)\in E}|[\Sigma^{-1}]_{ij}|\ \geq\ 4\kappa_{\Gamma}B\sqrt{\frac{\log d}{n}}$
holds, then $\mathrm{sign}(\widetilde{C}_{ij})=\mathrm{sign}([\Sigma^{-1}]_{ij})$ for all $(i,j)\in E$.
Theorem 3.4 establishes an error rate of order $\sqrt{\log d/n}$, support recovery, and sign consistency. While these rates align with classic results for the $\ell_{1}$-regularized estimation (Ravikumar et al., 2011), the conditions required to achieve them are different.
In the classic $\ell_{1}$-regularized precision matrix estimation, a uniform penalty is applied to all entries. The associated model selection consistency is established under the incoherence condition:
$\mu^{\ast}:=\max_{e\in S^{c}}\sum_{s\in S}|A_{es}|\leq 1-\alpha,\quad\alpha\in(0,1).$
This condition requires that no null edge is highly correlated with the set of true edges. Our perturbed estimation replaces this uniform requirement with the scale-adaptive incoherence condition, as shown in Assumption 3.3. The distinction lies in the penalty ratio $\lambda^{*}_{s}/\lambda^{*}_{e}$, which essentially compares the marginal standard deviation of the variables involved in a true edge versus a null edge.
We claim that the scale-adaptive incoherence condition embedded in our perturbed estimation (10) is superior over the classic one in the $\ell_{1}$-regularized estimation under heteroscedastic settings. This can be seen from the following proposition.
Proposition 3.5.
Suppose the covariance matrix $\Sigma$ admits the decomposition $\Sigma=DRD$, where $D=\mathrm{diag}(\sqrt{\Sigma_{11}},\dots,\sqrt{\Sigma_{dd}})$. Let $\Gamma^{\dagger}:=R\otimes R$. Define
$A^{\dagger}:=\Gamma_{S^{c}S}^{\dagger}(\Gamma_{SS}^{\dagger})^{-1}.$
If $10\delta<\sqrt{2(\pi)^{-1}}\sigma_{\min}$, then
$\mu^{\ast}=\max_{e\in S^{c}}\sum_{s\in S}|A_{es}|\leq\Big(\frac{\sigma_{\max}}{\sigma_{\min}}\Big)^{2}\|A^{\dagger}\|_{\infty},\quad\psi^{\ast}=\max_{e\in S^{c}}\sum_{s\in S}|A_{es}|\frac{\lambda^{\ast}_{s}}{\lambda^{\ast}_{e}}\leq\frac{21}{20}\Big(\frac{\sigma_{\max}}{\sigma_{\min}}\Big)\|A^{\dagger}\|_{\infty}.$
Compared to the standard incoherence $\mu^{\ast}$, the scale-adaptive version $\psi^{\ast}$ reduces the incoherence upper bound by a factor of $\sigma_{\max}(\sigma_{\min})^{-1}$, indicating a possible relaxation of the incoherence condition under the heteroscedastic settings. The improvement will be more significant as the difference between the maximum marginal standard deviation and the minimum one becomes larger.
@@ 4 Regularization Effect
We characterize this regularization effect in the following proposition.
@@
Suppose the random variable $\bm{x}$ follows a distribution $\bm{P}$ with $\mathbb{E}_{\bm{P}}[\|\bm{x}\|_{p}]<\infty$ and $\bm{P}([C\bm{x}]_{i}=0)=0,i=1,\ldots,d$, $C$ is a positive definite matrix with $\|C\|_{p\to q}<\infty$, and $p\in(1,\infty]$ with $1/p+1/q=1$. Then, we have that
@@
$\bm{v}_{C\bm{x}}=\frac{\mathrm{sign}(C\bm{x})\odot|C\bm{x}|^{q-1}}{\|C\bm{x}\|_{q}^{q-1}}.$ (12)
@@
Since the term $\log\det C$ in the proposed framework (2) is deterministic, Proposition 4.1 focuses on the inner expectation in (2) to analyze the induced regularization under $\ell_{p}$-perturbation. Instead of only analyzing on the empirical distribution, we consider a more general distribution $\bm{P}$ satisfying certain regularity conditions. The stated regularity conditions in Proposition 4.1 ensure well-posedness and integrability.
@@
$=$ $\mathbb{E}_{\bm{P}}[\bm{x}^{\top}C\bm{x}]+2\delta\mathbb{E}_{\bm{P}}\left[\|C\bm{x}\|_{1}\right]+\delta^{2}\mathbb{E}_{\bm{P}}\left[\mathrm{sign}(C\bm{x})^{\top}C\mathrm{sign}(C\bm{x})\right]+o(\delta^{2}),$
@@
The first-order regularization term arises from Hölder’s inequality and penalizes the $\ell_{q}$ norm of the gradient direction, where $1/p+1/q=1$. The vector $\bm{v}_{C\bm{x}}$ in the second-order term is the $\ell_{p}$-unit vector that achieves the associated Hölder dual pairing, as defined in (12). The second-order regularization term reflects the local curvature of the worst-case loss. Specifically, this term corresponds to the mean of the Rayleigh quotient of $C$ evaluated with respect to the vector $C\bm{x}$ (when $p=2$) or $\mathrm{sign}(C\bm{x})$ (when $p=\infty$). If $C$ has the eigen-decomposition $C=\sum_{j}\lambda_{j}\bm{u}_{j}\bm{u}_{j}^{\top}$, the associated Rayleigh quotient represents a weighted average of the eigenvalues. For $p=2$, it takes the form:
@@
$\mathrm{sign}(C\bm{x})^{\top}C\mathrm{sign}(C\bm{x})=\sum_{j}\lambda_{j}(\bm{u}_{j}^{\top}\mathrm{sign}(C\bm{x}))^{2}.$
@@ 5 Asymptotic Distribution
$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}(\bm{x}_{i}\bm{x}_{j},\bm{x}_{i^{\prime}}\bm{x}_{j^{\prime}}).$
@@
Firstly, the proposed perturbed estimation $\widehat{C}$ is a consistent estimator of the precision matrix $\Sigma^{-1}$, i.e., $\widehat{C}\to_{p}\Sigma^{-1}$, when the perturbation magnitude shrinks as the sample size increases. Then, the proposed perturbed estimation has a convergence order of $\sqrt{n}$ when $\gamma\geq 1/2$ and a slower convergence order of $n^{\gamma}$ when $0<\gamma<1/2$.
@@
$\widetilde{C}\in\arg\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\omega_{k}|C_{kj}|+\delta^{2}\|C\|_{1,1}\right\}.$ (15)
@@
    $+\sqrt{\frac{8}{\pi}}\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\Sigma_{kk}\left(\bm{U}_{jk}\mathrm{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|\bm{U}_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
@@
    $+\sqrt{\frac{8}{\pi}}\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\Sigma_{kk}\left(\bm{U}_{jk}\mathrm{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|\bm{U}_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
@@
$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}(\bm{x}_{i}\bm{x}_{j},\bm{x}_{i^{\prime}}\bm{x}_{j^{\prime}}).$
Corollary 5.3 (Sparsity Recovery).
Let $\widetilde{C}$ be the minimizer of problem (9). Suppose that $\delta=\eta n^{-1/2}$ with $\eta>0$. Then, as $n\to\infty$, the asymptotic distribution of $\widetilde{C}_{ij}$ has a positive probability mass at 0 whenever $\Sigma^{-1}]_{ij}=0$.
@@
Corollary 5.3 follows immediately from Theorem 5.2 and Proposition 3.8 in Xie and Huo (2025).
Theorem 5.2 shows that the estimation $\widetilde{C}$ in problem (15) shares the same asymptotic normal distribution of the estimation $\widehat{C}$ in problem (14). However, for $\gamma=1/2$, the asymptotic distribution of $\widetilde{C}$ is non-normal. The term $\Sigma_{kk}$ presented in Theorem 5.2 is resulted from the adversarial perturbation in the framework (15), showing the asymptotic effect of the scale-adaptivity and adversarial robustness discussed in Section 3.1 and Section 3.2.
In addition, our asymptotic analysis demonstrates that the sparsity recovery can be achievable when $\gamma=1/2$, seeing Corollary 5.3. Recall the high dimensional finite-sample analysis implies that model selection consistency is obtained when $\delta$ is in the order of $n^{-1/2}$. This provides a unification of finite-sample and asymptotic perspectives for $\ell_{\infty}$-perturbed estimation, demonstrating that both regimes demand an equivalent scaling of $\delta\asymp n^{-1/2}$ to guarantee model selection consistency.
@@
$p\in(1,\infty]$, problem (2) $p=\infty$, surrogate, problem (9) Rate
@@ 6 Numerical Experiments
In this section, we will deliver numerical experiments for the proposed adversarially perturbed precision matrix estimation scheme and compare it with the $\ell_{1}$-regularized precision matrix estimation (Friedman et al., 2008). The experiments are implemented by R. The code url:https://github.com/YilingXie27/Perturbed-Precision-Matrix-Estimation.
We will focus on the surrogate $\ell_{\infty}$-perturbed matrix estimation framework (9) discussed in Section 3. While the $\ell_{\infty}$-perturbed precision matrix estimation has penalties dependent on the observed samples, the $\ell_{1}$-regularized precision matrix estimation (Friedman et al., 2008) has a constant penalty, as shown in the following:
@@
We may use the following metrics to evaluate the numerical performance : true negative rate (TNR), true positive rate (TPR), accuracy (ACC), and Matthews Correlation Coefficient (MCC) (Guilford, 1954). The metrics are defined as follows:
@@
$\mathrm{ACC}=\frac{\mathrm{TN}+\mathrm{TP}}{\mathrm{TN}+\mathrm{FP}+\mathrm{TP}+\mathrm{FN}},$
@@
6.1 Experiments with Synthetic Data
We will carry out the numerical experiments on the following synthetic models.
Model 1. An $AR(2)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.5$ and $c_{i,i-2}=c_{i-2,i}=0.25$, for $i=1,\ldots,d$.
Model 2. An $AR(3)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.4$ and $c_{i,i-2}=c_{i-2,i}=c_{i,i-3}=c_{i-3,i}=0.2,$ for $i=1,\ldots,d$.
Model 3. An $AR(4)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.4$, $c_{i,i-2}=c_{i-2,i}=c_{i,i-3}=c_{i-3,i}=0.2$ and $c_{i,i-4}=c_{i-4,i}=0.1,$ for $i=1,\ldots,d$.
Model 4. Star model with every node connected to the first node, with $c_{i,i}=1$, $c_{1,i}=c_{i,1}=0.2$ and $c_{i,j}=0$ otherwise, for $i,j=1,\ldots,d$.
Model 5. Circle model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.5$ and $c_{1,d}=c_{d,1}=0.4$, for $i=1,\ldots,d$.
We inject heteroskedastic marginal variances by applying a diagonal rescaling into the models above to illustrate the superiority of the perturbed method under heterogeneous variance settings, as discussed in Section 3.2. Define a diagonal scaling matrix $S=\mathrm{diag}(s_{1},\ldots,s_{d})$ with $s_{i}=10$ for $i=1,\ldots,5$ and $s_{i}=1$ for $i=6,\ldots,d$. We then set the data-generating covariance to $\Sigma_{S}=S\Sigma S,$. Because $D$ is diagonal, the sparsity pattern of the corresponding precision matrix is unchanged, while the first five variables have substantially larger scales.